Consider the following short time series data set: t1234 5 6789101112 13 14 15 16 17 1819 20 x13 25 20 2 4 12 8 7 19 71 20 5 12 18 12 20 85 Evaluate each of the following expressions (i.e., give a numerical answer): a. ▽x13 c. ▽2x13 d. Vx13 e. 1-B1)x1 f. (1-0.7B +0.5B2)1-B)x13
Answer #1

Solution :

We will consider the given short time series data set where "t" denotes the time period and "xt" denotes the observation at the time period "t". We have to evaluate the given expressions!

We know the formulae for the given expressions. They are given below :

▽ is known as the Backward Difference Operator

where, riangledown X_t=X_t-X_{t-1} & riangledown^2X_t= riangledown( riangledown X_t)

B is known as the Backward Shift Operator

where. BXt = Xt-1 n general, BnXt = Xt-m 1 m > 0

A is known as the Forward Difference Operator

where, riangle X_t=X_{t+1}-X_{t} & riangle^2X_t= riangle( riangle X_t)

________________________________________________________________________________________________

(a) To find : Tx13

riangledown x_{13}=x_{13}-x_{13-1}=x_{13}-x_{12}=20-1=19

Thus, we have, Vx13 19 Ans)

__________________________

mathbf{(b) To find : B^2 x_{13}}

B^2x_{13}=B(Bx_{13})=B(x_{12})=x_{11}=7

2 Thus, we have, Bx137 (Ans)

__________________________

(c) To find : Д2x13

riangle^2 x_{13}= riangle( riangle x_{13})= riangle(x_{14}-x_{13})= riangle x_{14}- riangle x_{13}=(x_{15}-x_{14})-(x_{14}-x_{13})

  =x_{15}-x_{14}-x_{14}+x_{13}=x_{15}-2*x_{14}+x_{13}=12-2*5+20=12-10+20=22

mathbf{Thus, we have, riangle^2x_{13}=22.......................(Ans)}

__________________________

(d) To find : V2x13

riangledown^2 x_{13}= riangledown( riangledown x_{13})= riangledown(x_{13}-x_{12})= riangledown x_{13}- riangledown x_{12}=(x_{13}-x_{12})-(x_{12}-x_{11})

  =x_{13}-x_{12}-x_{12}+x_{11}=x_{13}-2*x_{12}+x_{11}=20-2*1+7=25

mathbf{Thus, we have, riangledown^2 x_{13}=25.......................(Ans)}

__________________________

mathbf{(e) To find : (1-B^{12}) x_{13}}

(1-B^{12}) x_{13}=x_{13}-B^{12}x_{13}=x_{13}-x_{13-12}=x_{13}-x_1=20-13=7

mathbf{Thus, we have, (1-B^{12}) x_{13}=7.......................(Ans)}

__________________________

(f) To find (1- 0.7B +0.5B2) (1 - B)xi3

(1-0.7B+0.5B^2)(1-B) x_{13}=[(1-0.7B+0.5B^2)-B(1-0.7B+0.5B^2)]x_{13}

  =(1-0.7B+0.5B^2-B+0.7B^2-0.5B^3)x_{13}

  =x_{13}-0.7Bx_{13}+0.5B^2x_{13}-Bx_{13}+0.7B^2x_{13}-0.5B^3x_{13}

  =x_{13}-0.7x_{13-1}+0.5x_{13-2}-x_{13-1}+0.7x_{13-2}-0.5x_{13-3}

  =x_{13}-0.7x_{12}+0.5x_{11}-x_{12}+0.7x_{11}-0.5x_{10}

  =20-0.7*1+0.5*7-1+0.7*7-0.5*19

  20-0.7+ 3.5-1 + 4.9-9.5 = 17.2

mathbf{Thus, we have, (1-0.7B+0.5B^2)(1-B) x_{13}}=17.2.......................(Ans)}

__________________________